In his pocket Pat has 5 dimes, 7 nickels, and 4 pennies. He selects 4 coins.
a.)How many ways can he select 4 dimes?
b.)How many ways can he select 4 pennies?
c.) How many ways can he select 3 nickels and one dime?
d.) How many ways can he select exactly 3 nickels?
i tried to use the combination of all the possibilities and do that for the first few (a, b, and d), but I wasn't sure if that was right. Really if I can just get the letter c I can probably get the rest of them okay.

Answer by stanbon(75887)

In his pocket Pat has 5 dimes, 7 nickels, and 4 pennies. He selects 4 coins.
a.)How many ways can he select 4 dimes?.........5C4 = 5
b.)How many ways can he select 4 pennies?.......4C4 = 1
c.) How many ways can he select 3 nickels and one dime?..7C3*5C1=35*5=175
d.) How many ways can he select exactly 3 nickels?....7C3=35
=============
Cheers,
Stan H.